# every label/.append style = font =

### v1

@@ III.6 Optimizing diode efficiencies: the role of the inhomogeneous field amplitude
Figure 13: Panel $(a)$ shows the critical currents $I_{c}^{+}$ (blue circles), $I_{c}^{-}$ (red circles), $I_{cv}^{+}$ (blue dots) and $I_{cv}^{-}$ (red dots) as a function of the current density in the side stripes $J$. Panel $(b)$ shows $\epsilon_{SDE}$ (blue circles) and $\epsilon_{VDE}$ (red circles) as a function of $J$. Top $x$ axis relates $J$ with the value of the inhomogeneous field at the sample edges $H_{e}$.
@@
In the above discussion, we have studied the efficiency of our diode system considering a fixed antisymmetric field profile, given by $J_{1}=J_{2}=0.26$. Here we investigate how the diode efficiency changes when varying the amplitude of the external field profile. For that, we fix $J_{1}=J_{2}=J$ and vary the value of $J$. We perform the calculations for a linear current cycle with sweep time $\tau=10^{6}t_{GL}$ and $\eta=2\times 10^{-5}$. Panel $(a)$ of Fig. 13 shows the critical currents $I_{c}^{\pm}$ (blue and red circles) and $I_{cv}^{\pm}$ (blue and red dots). Both $I_{cv}^{+}$ and $I_{c}^{+}$ decrease monotonically with $J$, which follows from the fact that the inhomogeneous field profile adds to the self-field of the AC current in the positive current branch. Therefore, increasing $J$ contributes to increasing the number of vortices and antivortices as well as the rate of their collision, so that a lower current is required to initiate the dissipative regime. In contrast, $I_{cv}^{-}$ and $I_{c}^{-}$ present a non-monotonic dependence on $J$, which can be understood as follows. On one hand, increasing $J$ makes it necessary a larger value of the current magnitude to compensate the inhomogeneous field and induce the resistive state for negative current polarities. On the other, when $J$ becomes too strong, the inhomogeneous field depletes superconductivity near the edges so that lower values of the current magnitude are capable of driving the system into the normal state.
@@
This discrepancy in the behavior of the system for each polarity leads to the existence of an optimal value of $J$ that gives the largest efficiency. Panel $(b)$ of Fig. 13 shows $\epsilon_{SDE}$ (blue circles) and $\epsilon_{VDE}$ (red circles) as a function of $J$. As one can see, $J\approx 0.36$ provides the largest efficiency of the superconducting diode effect, while $J\approx 0.40$ is the optimal value for the vortex diode effect.
@@ IV Conclusions
In conclusion, we have explored the interplay between the superconducting diode effect (SDE) and the vortex diode effect (VDE) in thin superconducting films subjected to external asymmetric flux profiles. Our findings highlight the potential of inhomogeneous fields to create highly asymmetric conditions for vortex entrance and exit, unveiling highly efficient vortex diode effects in plain superconducting films. Notably, the diode effect observed in our study does not require sophisticated nanofabrication techniques, as the inhomogeneous field can be induced by plain ferromagnetic films underneath the superconductor or by currents injected in microsized superconducting wires coplanar with the main superconducting film.
Moreover, our results demonstrate that the perfect antisymmetric flux profile represents the most efficient scenario for the diode effect. This profile maximizes the difference between the Bean-Livingstone barriers in opposite edges, leading to the highest vortex diode efficiency. Additionally, the built-in coexistence of vortices and antivortices in this scenario promotes a high rate of vortex-antivortex annihilations at the equatorial line of the sample during the positive current branch. This induces hot spots, even in the early stages of the flux-flow regime, bringing forward the transition to the normal state. In contrast, during the negative current branch, the current self-field compensates the external antisymetric field profile, eliminating vortices and antivortices from the sample, thus inducing the system into a perfectly superconducting state. Overall, for a wide range of parameters, the system behaves as a perfect superconducting half-wave rectifier, presenting no resistance or heat flow for half the current cycle.
@@ Acknowledgements.
LRC and ES thank the Brazilian Agency FAPESP for financial support, grant numbers 20/03947-2 and 20/10058-0, respectively. CCSS thanks financial support from Conselho Nacional de Desenvolvimento Científico e Tecnológico - Brasil (CNPq), grant number 312240/2021-0, and Universidade Federal de Pernambuco, Edital Produtividade Qualificada. This work was financed in part by Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES), Finance Code 001.

### v2

@@ III.6 Optimizing diode efficiencies: the role of the inhomogeneous field amplitude and shape
Figure 13: Panel $(a)$ shows the critical currents $I_{c}^{+}$ (blue circles), $I_{c}^{-}$ (red circles), $I_{cv}^{+}$ (blue dots) and $I_{cv}^{-}$ (red dots) as a function of the current density in the side stripes $J$. Panel $(b)$ shows $\epsilon_{NSD}$ (blue circles) and $\epsilon_{FFD}$ (red circles) as a function of $J$. Top $x$ axis relates $J$ with the value of the inhomogeneous field at the sample edges $H_{e}$.
@@
In the above discussion, we have studied the efficiency of our diode system considering a fixed antisymmetric field profile, given by $J_{1}=J_{2}=0.26$ and distance $s=5\xi(0)$ between the superconducting film and the lateral stripes. Here we investigate how the diode efficiency changes when varying the amplitude and shape of the external field profile. For that, we fix We start by maintaining $s=5\xi(0)$, fixing $J_{1}=J_{2}=J$ and vary varying the value of $J$. We perform the calculations for a linear current cycle with sweep time $\tau=10^{6}t_{GL}$ and $\eta=2\times 10^{-5}$. Panel $(a)$ of Fig. 13 shows the critical currents $I_{c}^{\pm}$ (blue and red circles) and $I_{cv}^{\pm}$ (blue and red dots). Both $I_{cv}^{+}$ and $I_{c}^{+}$ decrease monotonically with $J$, which follows from the fact that the inhomogeneous field profile adds to the self-field of the AC current in the positive current branch. Therefore, increasing $J$ contributes to increasing the number of vortices and antivortices as well as the rate of their collision, so that a lower current is required to initiate the dissipative regime. In contrast, $I_{cv}^{-}$ and $I_{c}^{-}$ present a non-monotonic dependence on $J$, which can be understood as follows. On one hand, increasing $J$ makes it necessary a larger value of the current magnitude to compensate the inhomogeneous field and induce the resistive state for negative current polarities. On the other, when $J$ becomes too strong, the inhomogeneous field depletes superconductivity near the edges so that lower values of the current magnitude are capable of driving the system into the normal state.
@@
This discrepancy in the behavior of the system for each polarity leads to the existence of an optimal value of $J$ that gives the largest efficiency. Panel $(b)$ of Fig. 13 shows $\epsilon_{NSD}$ (blue circles) and $\epsilon_{FFD}$ (red circles) as a function of $J$. As one can see, $J\approx 0.36$ provides the largest efficiency of the superconducting diode effect, while $J\approx 0.40$ is the optimal value for the vortex diode effect.
Figure 14: Panel $(a)$ shows the diode efficiency as a function of the separation between the superconducting film and the lateral wires carrying a current chosen as to induce $H_{e}=0.69H_{c2}(T)$. Panel $(b)$ presents the inhomogenoeus field profile for different values of $s/L_{y}$ (solid lines) and the self-field profile of the central superconducting film (black dashed line). All curves are normalized by the field at the edges
@@
Let us now investigate how the diode efficiency depends on the shape of the inhomogeneous field. Different shapes can be induced by changing the separation $s$ between the wires and the main superconducting film: for small separations, the profile is more concentrated at the film edges, while for large separations the profile tends to a linear shape (see solid curves in panel $(b)$ of Fig. 14). In particular, we note that for $s/L_{y}=0.025$ the profile resembles the self-field induced by the applied current in the absence of vortices. In panel $(a)$ of Fig. 14, we show the diode efficiencies as a function of $s/L_{y}$. For each value of $s$, the current density in the wires, $J$, was chosen so as to induce $H_{e}=0.69H_{c2}(T)$ at the film edges. It turns out that the optimum efficiency is provided precisely by $s/L_{y}=0.025$. This leads to the conclusion that the most efficient way of promoting the superconducting diode effect is by using inhomogeneous field textures that replicate the self-field so as to fully compensate the effect of the applied current in a given polarity, thus keeping the device in the superconducting state.
@@ IV Conclusions
In conclusion, we have explored the interplay between different operation modes of the superconducting diode effect in thin superconducting films subjected to external asymmetric flux profiles. Our findings highlight the potential of inhomogeneous fields to create highly asymmetric conditions for vortex entrance and exit, unveiling highly efficient vortex diode effects in plain superconducting films. Notably, the diode effect observed in our study does not require sophisticated nanofabrication techniques, as the inhomogeneous field can be induced by plain ferromagnetic films underneath the superconductor or by currents injected in microsized superconducting wires coplanar with the main superconducting film. A centimeter-sized version of the latter setup, but using an in-plane homogeneous field, was explored long ago in [52], finding diode efficiencies reaching up to 67%, similar to ours. However, it is unclear whether the origin of non-reciprocity is the same in both studies, since in our proposed micron-scale setup no external homogeneous field is required.
Moreover, our results demonstrate that the perfect antisymmetric flux profile represents the most efficient scenario for the diode effect. This profile maximizes the difference between the Bean-Livingstone barriers in opposite edges, leading to the highest vortex diode efficiency. Additionally, the built-in coexistence of vortices and antivortices in this scenario promotes a high rate of vortex-antivortex annihilations at the equatorial line of the sample during the positive current branch. This induces hot spots, even in the early stages of the flux-flow regime, bringing forward the transition to the normal state. In contrast, during the negative current branch, the current self-field compensates the external antisymetric field profile, eliminating vortices and antivortices from the sample, thus inducing the system into a perfectly superconducting state. Overall, for a wide range of parameters, the system behaves as a superconducting half-wave rectifier, presenting no resistance or heat flow for half the current cycle.
@@ Acknowledgements.
We thank A. V. Silhanek for bringing Reference [52] to our attention. LRC and ES thank the Brazilian Agency FAPESP for financial support, grant numbers 20/03947-2 and 20/10058-0, respectively. CCSS thanks financial support from Conselho Nacional de Desenvolvimento Científico e Tecnológico - Brasil (CNPq), grant number 312240/2021-0, and Universidade Federal de Pernambuco, Edital Produtividade Qualificada. This work was financed in part by Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES), Finance Code 001.